OpenStudy (anonymous):

Please help! I'm stuck. Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). I've done: y-3 = 2(x+6) y+3 = 2x - 12 +3 + 3 y = 2x - 9 I'm stuck now, can someone please work it out for me:/

OpenStudy (anonymous):

Perpendicular, slope would be -1/2

OpenStudy (anonymous):

There asking for the equation, and the possible answers are y = 2x + 6 y = −one halfx + 3 y = −one halfx + 6 y = 2x + 3

OpenStudy (anonymous):

So what you'll need to do is set up an equation to find what y will equal when x=0

OpenStudy (anonymous):

y=-1/2x+6

OpenStudy (anonymous):

You can check by plotting your point at (6,3) since you know slope is -1/2, you would move from (6,3) up one unit, to the left two units until you cross the y axis.. where you cross the y axis is your y-intercept.. or "b"

OpenStudy (anonymous):

line is perpendicular to the given so the slope is -ve reciprocal i.e -1/2 and a point is given, solve for the equation of line you will get [y-3]=(-1/2)[x-6] ==> y=-x/2 +6